# Topological zero modes and bounded modes at smooth domain walls: exact solutions and dualities

### v1

@@ Introduction
$\begin{cases}v_{L,R}^{2}+m_{L,R}\geq 0\rightarrow\text{exp.~{}decay}\\ v_{L,R}^{2}+m_{L,R}<0\rightarrow\text{exp.~{}damped oscill.}\end{cases}\text{ for }x\to\mp\infty$ (8)
@@
Table 2: The number of modes and pseudospin $s$ satisfying the boundary conditions $\varphi(\pm\infty)=0$ on the infinite interval $-\infty<x<\infty$ and $\varphi(0)=\varphi(\infty)=0$ on the semi-infinite interval $0\leq x<\infty$ depending on the asymptotic values $m_{L,R},v_{L,R}$, and $W_{L,R}$. The modes show exponential decay if ${v}_{L}^{2}+m_{L}\geq 0$ or exponentially damped oscillations if ${v}_{L}^{2}+m_{L}<0$. The asymptotic behavior of the solutions is completely determined by the decay rates and decay lengths $\mu_{L,R}^{\pm}=1/\xi_{L,R}^{\pm}=\mathrm{Re}(\alpha_{L,R}^{\pm})$ and the momenta and wavelengths $\kappa_{L,R}=2\pi/\lambda_{L,R}=|\mathrm{Im}(\alpha_{L,R}^{\pm})|$ responsible for oscillatory behaviors. Furthermore, $\kappa_{L,R}=0$ for ${v}_{L,R}^{2}+m_{L,R}\geq 0$, with $\mu_{L,R}^{+}=\mu_{L,R}^{-}$ for ${v}_{L,R}^{2}+m_{L,R}=0$, while $\mu_{L,R}^{+}=\mu_{L,R}^{-}=\mp sv_{L,R}$ and $\kappa_{L,R}\neq 0$ for ${v}_{L,R}^{2}+m_{L,R}<0$.
@@
On the contrary, zero modes localized at a smooth domain wall have "hair" in the sense that they are nonfeatureless, i.e., cannot be described entirely by a small set of numbers, at least not at short length scales. Indeed, in the case $w<\xi,\lambda$, modes localized at a smooth domain wall appear featureless at large distances $|x|>\xi,\lambda>w$ but not at short distances $|x|<w$. As long as they are looked at from a large distance $|x|>\xi,\lambda>w$, these modes are effectively hairless, i.e., they can be treated as effectively featureless modes at long length scales, being as good as zero modes localized at sharp phase boundaries, and being characterized only by a set of few numbers ($\xi,\lambda$ or $\alpha$). We refer to this case as zero modes with "short hair". Dramatically different is the case where $w>\xi,\lambda$. In this case, the features of the wavefunctions cannot be neglected since the length scales associated with these features are larger than the decaying lengths of the modes. Hence, these wavefunctions are nonfeatureless at all length scales, i.e., cannot be described only by a set of few numbers $\xi,\lambda$ or $\alpha$ at any scale. We refer to this case as zero modes with "long hair". Another way to look at the difference between the case $w<\xi,\lambda$ and the case $w>\xi,\lambda$ is to say that the spatial extension of the mode is larger than the spatial extension of the smooth domain wall. Notice also that the property of being a short-hair $w<\xi,\lambda$ or long-hair $w>\xi,\lambda$ mode depends on whether we look at the mode from the left side (i.e., $x<0$) or from the right side (i.e., $x>0$ for a smooth domain wall at $x=0$). We notice that, even though their topological nature is not apparent at all length scales, the existence of the short-hair $w<\xi,\lambda$ and long-hair $w>\xi,\lambda$ modes at smooth domain walls is still a direct consequence of the bulk-boundary correspondence. In the context of one-dimensional topological superconductors, featureless, hairless zero modes correspond to textbook examples of Majorana zero modes localized at a sharp boundary between topologically nontrivial phase and a trivial phase (e.g., the vacuum), whereas nonfeatureless zero modes in the short-hair or long-hair regimes correspond to different incarnations of quasi-Majorana modes which appear at smooth boundaries[23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
@@
Figure 1: Zero modes on the interval $-\infty<x<\infty$ with a smooth domain wall described by the fields $m(x)$ and $v(x)$ with a width smaller than or comparable with the characteristic decay lengths $w\lesssim\xi$. Upper panels show the wavefunction, middle panels show the logarithm of the absolute value of the wavefunction, and lower panes show the fields $v(x)$, $m(x)$, and $v(x)^{2}+m(x)$. Different columns correspond to different choices of the parameters $m_{0,1,2}$ and $v_{1,2}$: (a) single mode with exponential decay on both sides of a smooth domain wall with S-shaped field $m(x)$ and constant $v(x)$; (b) single mode with exponentially damped oscillations on the right side and exponential decay on the left side of a smooth domain wall with S-shaped fields $m(x)$ and $v(x)$; (c) two modes with exponentially damped oscillations on both sides of a smooth domain wall with S-shaped field $v(x)$ and constant $m(x)$ (we consider the two linear combinations $\varphi_{1}\pm\varphi_{2}$ giving real wavefunctions); (d) single mode with exponential decay on both sides of a smooth domain wall with asymmetric Pöschl–Teller field $m(x)$ and constant $v(x)$; (e) single mode with exponentially damped oscillations on the right side and exponential decay on the left side of a smooth domain wall with asymmetric Pöschl–Teller field $m(x)$ and S-shaped field $v(x)$; (f) two modes with exponentially damped oscillations on both sides of a smooth domain wall with symmetric Pöschl–Teller field $m(x)$ and S-shaped field $v(x)$ (linear combinations $\varphi_{1}\pm\varphi_{2}$).
@@
The analytic solutions we found describe zero modes at smooth domain walls in all regimes, i.e., featureless and nonfeatureless cases with short-hair $w<\xi$ and long-hair $w>\xi$. Moreover, the long-distance asymptotic behavior for $x\gg w$ is completely determined by the exponents $\mu$ and $\kappa$, and by the signs of the quantities $m_{L,R}$ and $v_{L,R}^{2}+m_{L,R}$. However, to gain a physical intuition on the localization of the zero modes at short distances $x\lesssim w$, one can analize the spatial dependence of the quantities $m(x)$ and $v(x)^{2}+m(x)$ and define the local topological invariant
@@
Hence, the local topological invariant encodes the information about the localization of the zero modes, which may not localize near the center of the smooth domain wall if $w\gg\xi$. On the other hand, we have seen that the quantity $v_{L,R}^{2}+m_{L,R}$ determines whether the zero modes exhibit exponentially damped oscillations ($v_{L,R}^{2}+m_{L,R}<0$) or a simple exponential decay without oscillations ($v_{L,R}^{2}+m_{L,R}\geq 0$). Similarly, one can surmise that the sign of the function $v(x)^{2}+m(x)$ will indicate the presence of oscillations for $v(x)^{2}+m(x)<0$ and the absence of oscillations for $v(x)^{2}+m(x)\geq 0$ at any given point $x$. In particular, it is possible that the rate of the oscillations may change considerably around the center of the domain wall $|x|\lesssim w$. More remarkably, oscillations may appear only within an interval $|x|\lesssim x_{\text{osc}}$ if $v(x)^{2}+m(x)<0$ for $|x|<x_{\text{osc}}$ and $v(x)^{2}+m(x)=v_{L,R}^{2}+m_{L,R}\geq 0$ for $|x|>x_{\text{osc}}$ or conversely, only for $|x|>x_{\text{osc}}$ if $v(x)^{2}+m(x)\geq 0$ for $|x|<x_{\text{osc}}$ and $v(x)^{2}+m(x)=v_{L,R}^{2}+m_{L,R}<0$ for $|x|>x_{\text{osc}}$.
@@
Figure 2 show zero modes on the interval $-\infty<x<\infty$ in the case where the width of the smooth domain wall $w$ is smaller than or comparable with the characteristic decay lengths $\xi$, for different choices of the parameters $m_{0,1,2}$ and $v_{1,2}$, plotted using Eqs. 4 and 5. These figures show examples of modes with exponential decay or exponentially damped oscillations. The exponential decay of the modes is clearly visible, with the exception of a small region for $|x|<w<\xi$: the modes have "short-hair", i.e., they are nonfeatureless for $|x|\lesssim w$ but can be regarded as featureless for distances $|x|\gtrsim w$. Figure 2 show cases analogous to Fig. 2, but where the width of the smooth domain wall $w$ is larger than the characteristic decay lengths $\xi$. The most noticeable difference with the previous case is that the exponential decay of the modes is now visible only at large distances $|x|\gtrsim w>\xi$: the modes have "long-hair", i.e., they are nonfeatureless at finite length scales.
The local features of the zero modes become prominent when $w\gg\xi$ (long-hair) and when the field around the domain wall is large. This is particularly visible in Fig. 2(d) where $w>\xi$, with oscillations visible in a wide region close to the center where $v(x)^{2}+m(x)<0$ but disappear at larger distances where $v(x)^{2}+m(x)>0$, and to some extent in Fig. 2(d) where $w<\xi$, with oscillations visible in a narrower region. Conversely, in Fig. 2(f), oscillations seem to occur only at large distances where $v(x)^{2}+m(x)<0$ but disappear at smaller distances where $v(x)^{2}+m(x)>0$. Even more remarkably, the zero mode in Fig. 2(d) is localized far away from the center of the domain wall in correspondence of a point $|x_{1}^{*}|>\xi$, while the two zero modes in Fig. 2(f) are localized around the two points $|x_{1,2}^{*}|>\xi$ where the local topological invariant $W(x)$ changes and the field $m(x)$ changes its sign.
@@
$\widetilde{\varphi}(x)=(2\cosh(x/(2w))^{swv_{1}}\mathrm{e}^{s(2v_{0}+v_{1})x/2}\varphi(x),$ (19)
@@
$\widetilde{\varphi}_{1,2}(x)=$ $\frac{\mathrm{e}^{\pm({q_{L}}+{q_{R}})x/2}}{\left(2\cosh{(x/2w)}\right)^{{\pm w(q_{L}}-{q_{R}})}}F_{1,2}(x),$ (21)
@@
1) general case $\widetilde{\varphi}_{1,2}(x)=\dfrac{\mathrm{e}^{\pm\left(q_{L}+q_{R}\right)x/2}}{\left(2\cosh{(x/2w)}\right)^{\pm w\left(q_{L}-q_{R}\right)}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
Notice that we required different kinds of boundary conditions for the zero-energy solutions of the generalized Jackiw-Rebbi equation and for the associated Schrödinger equation. In the case of the generalized Jackiw-Rebbi equation, we are concerned about normalizable modes that satisfy the Dirichlet boundary conditions $\varphi(\pm\infty)=0$. These normalizable modes do not exist, in general, for the associated Schrödinger equation. In this case, we only require that the modes are bounded $|\widetilde{\varphi}(x)|<\infty$ for $x\to\pm\infty$. These modes behave essentially as plane waves at long distances or decay exponentially.
@@
The bounded modes $\widetilde{\varphi}_{1,2}(x)$ on the infinite interval $-\infty<x<\infty$ with $\widetilde{W}_{L}=\widetilde{W}_{R}=1$, i.e., $E\geq V_{L,R}$, describe the transmission and reflection of particles with energies greater than the height of the domain wall $E\geq\max(V_{L,R})$. These modes are wavelike modes with oscillatory having momentum with momenta $\kappa_{L,R}=\mathrm{Im}(q_{L,R})=\sqrt{E-V_{L,R}}$, respectively on each side of the domain wall. Hence, these modes are completely delocalized. When $w\to 0$ and $V(x)=V_{L}=V_{R}$, this simply reduces to the case of a free particle with momentum $\kappa=\sqrt{E-V}$.
@@
Moreover, the bounded mode on the semi-infinite interval $0\leq x<\infty$ $\widetilde{W}_{R}=1$, i.e., $E\geq V_{R}$ describes the reflection of a particle with momentum $\kappa=\sqrt{E-V_{R}}$ off an infinite wall with smooth spatial variations of the potential near $x=0$, which for $w\to 0$ reduces to the well-known case describing the reflection of a free particle scattering off an infinite wall.
@@
The bounded mode $\widetilde{\varphi}_{1}(x)$ with $\widetilde{W}_{L}=0$ and $\widetilde{W}_{R}=1$ with $V_{L}>E\geq V_{R}$ correspond via the duality transformation to the zero mode $\varphi_{1}^{s}(x)$ with $W_{L}=0$ and $|W_{R}|=1$ if we assume that $|m_{L}|>v_{L}^{2}$ and $v_{R}\neq 0$. On the right side, the bounded mode is oscillating and the zero node is exponentially decaying with damped oscillations and decay rate $|v_{R}|$. The oscillations have same momentum $\kappa_{R}=\mathrm{Im}(\sqrt{v_{R}^{2}+m_{R}})=\mathrm{Im}(\sqrt{V_{R}-E})$. Again, in the limiting cases where this momentum vanishes (i.e., if $E=V_{R}$) the bounded mode converges to a constant and the zero mode decay exponentially without oscillations. On the left side, the bounded mode and zero mode decay exponentially with decay rates $\mu_{L}=q_{L}$ and $\mu_{L}^{+}=-sv_{L}\pm q_{L}$, respectively. Same considerations can be done for the bounded mode $\widetilde{\varphi}_{2}(x)$ and zero mode $\varphi_{2}^{s}(x)$.
Finally, the bounded mode on the semi-infinite interval $0\leq x<\infty$ with $\widetilde{W}_{R}=1$, i.e., $E\geq V_{R}$, correspond via the duality transformation to the zero mode if we assume $v_{R}\neq 0$. The bounded mode is oscillating and the zero node is exponentially decaying with damped oscillations, with same momentum $\kappa_{R}=\mathrm{Im}(\sqrt{v_{R}^{2}+m_{R}})=\mathrm{Im}(\sqrt{V_{R}-E})$. In the limiting case, $E=V_{R}$) the bounded mode converges to a constant and the zero node decay exponentially without oscillations. The duality between the bounded modes and the zero modes is summarized in Table 5.
@@
[…]
@@
[…]

### v2

@@ Introduction
$\begin{cases}v_{L,R}^{2}+m_{L,R}\geq 0\rightarrow\text{exp.~{}decay}\\ v_{L,R}^{2}+m_{L,R}<0\rightarrow\text{exp.~{}damped oscill.}\end{cases}\text{ for }x\to\mp\infty.$ (10)
@@
Table 2: The number of modes and pseudospin $s$ satisfying the boundary conditions $\varphi(\pm\infty)=0$ on the infinite interval $-\infty<x<\infty$ and $\varphi(0)=\varphi(\infty)=0$ on the semi-infinite interval $0\leq x<\infty$ depending on the asymptotic values $m_{L,R},v_{L,R}$, and $W_{L,R}$. The modes show exponential decay if ${v}_{L}^{2}+m_{L}\geq 0$ or exponentially damped oscillations if ${v}_{L}^{2}+m_{L}<0$. The asymptotic behavior of the solutions is completely determined by the decay rates and decay lengths $\mu_{L,R}^{\pm}=1/\xi_{L,R}^{\pm}=\mathrm{Re}(\alpha_{L,R}^{\pm})$ and the momenta and wavelengths $\kappa_{L,R}=2\pi/\lambda_{L,R}=|\mathrm{Im}(\alpha_{L,R}^{\pm})|$ responsible for oscillatory behaviors. Furthermore, ${v}_{L,R}^{2}+m_{L,R}\geq 0$ yields $\kappa_{L,R}=0$, and the limiting case ${v}_{L,R}^{2}+m_{L,R}=0$ yields $\mu_{L,R}^{+}=\mu_{L,R}^{-}$ and $\kappa_{L,R}=0$. Also, ${v}_{L,R}^{2}+m_{L,R}<0$ yields $\mu_{L}^{+}=\mu_{L}^{-}=-sv_{L}$, $\mu_{R}^{+}=\mu_{R}^{-}=sv_{R}$, and $\kappa_{L,R}\neq 0$.
@@
On the contrary, zero modes localized at a smooth domain wall have "hair" in the sense that they are nonfeatureless, i.e., cannot be described entirely by a small set of numbers, at least not at short length scales. Indeed, in the case $w<\xi,\lambda$, modes localized at a smooth domain wall appear featureless at large distances $|x|>\xi,\lambda>w$ but not at short distances $|x|<w$. As long as they are looked at from a large distance $|x|>\xi,\lambda>w$, these modes are effectively hairless, i.e., they can be treated as effectively featureless modes at long length scales, being as good as zero modes localized at sharp phase boundaries, and being characterized only by a few numbers ($\xi,\lambda$, or $\alpha$). We refer to this case as zero modes with "short hair". Dramatically different is the case where $w>\xi,\lambda$. In this case, the features of the wavefunctions cannot be neglected since the length scales associated with these features are larger than the decaying lengths of the modes. Hence, these wavefunctions are nonfeatureless at all length scales, i.e., cannot be described only by a set of few numbers $\xi,\lambda$, or $\alpha$ at any scale. We refer to this case as zero modes with "long hair". Another way to look at the difference between the case $w<\xi,\lambda$ and the case $w>\xi,\lambda$ is to say that the spatial extension of the mode is larger than the spatial extension of the smooth domain wall. Notice also that the property of being a short-hair $w<\xi,\lambda$ or long-hair $w>\xi,\lambda$ mode depends on whether we look at the mode from the left side (i.e., $x<0$) or from the right side (i.e., $x>0$ for a smooth domain wall at $x=0$). We notice that, even though their topological nature is not apparent at all length scales, the existence of the short-hair $w<\xi,\lambda$ and long-hair $w>\xi,\lambda$ modes at smooth domain walls is still a direct consequence of the bulk-boundary correspondence. In the context of one-dimensional topological superconductors, featureless, hairless zero modes correspond to textbook examples of Majorana zero modes localized at a sharp boundary between a topologically nontrivial phase and a trivial phase (e.g., the vacuum), whereas nonfeatureless zero modes in the short-hair or long-hair regimes correspond to different incarnations of quasi-Majorana modes which appear at smooth boundaries[23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
@@
Figure 1: Zero modes on the interval $-\infty<x<\infty$ with a smooth domain wall described by the fields $m(x)$ and $v(x)$ with a width smaller than or comparable with the characteristic decay lengths $w\lesssim\xi$. Upper panels show the wavefunction, middle panels show the logarithm of the absolute value of the wavefunction, and lower panels show the fields $v(x)$, $m(x)$, and $v(x)^{2}+m(x)$. Different columns correspond to different choices of the parameters $m_{0,1,2}$ and $v_{1,2}$: (a) single mode with exponential decay on both sides of a smooth domain wall with S-shaped field $m(x)$ and constant $v(x)$; (b) single mode with exponentially damped oscillations on the right side and exponential decay on the left side of a smooth domain wall with S-shaped fields $m(x)$ and $v(x)$; (c) two modes with exponentially damped oscillations on both sides of a smooth domain wall with S-shaped field $v(x)$ and constant $m(x)$ (we consider the two linear combinations $\varphi_{1}\pm\varphi_{2}$ giving real wavefunctions); (d) single mode with exponential decay on both sides of a smooth domain wall with asymmetric Pöschl–Teller field $m(x)$ and constant $v(x)$; (e) single mode with exponentially damped oscillations on the right side and exponential decay on the left side of a smooth domain wall with asymmetric Pöschl–Teller field $m(x)$ and S-shaped field $v(x)$; (f) two modes with exponentially damped oscillations on both sides of a smooth domain wall with symmetric Pöschl–Teller field $m(x)$ and S-shaped field $v(x)$ (linear combinations $\varphi_{1}\pm\varphi_{2}$).
@@
The analytic solutions we found describe zero modes at smooth domain walls in all regimes, i.e., featureless and nonfeatureless cases with short-hair $w<\xi$ and long-hair $w>\xi$. Moreover, the long-distance asymptotic behavior for $x\gg w$ is completely determined by the exponents $\mu$ and $\kappa$, and by the signs of the quantities $m_{L,R}$ and $v_{L,R}^{2}+m_{L,R}$. However, to gain a physical intuition on the localization of the zero modes at short distances $x\lesssim w$, one can analyze the spatial dependence of the quantities $m(x)$ and $v(x)^{2}+m(x)$ and define the local topological invariant
@@
Hence, the local topological invariant encodes the information about the localization of the zero modes, which may not localize near the center of the smooth domain wall if $w\gg\xi$. On the other hand, we have seen that the quantity $v_{L,R}^{2}+m_{L,R}$ determines whether the zero modes exhibit exponentially damped oscillations ($v_{L,R}^{2}+m_{L,R}<0$) or a simple exponential decay without oscillations ($v_{L,R}^{2}+m_{L,R}\geq 0$). Similarly, one can surmise that the sign of the function $v(x)^{2}+m(x)$ will indicate the presence of oscillations for $v(x)^{2}+m(x)<0$ and the absence of oscillations for $v(x)^{2}+m(x)\geq 0$ at any given point $x$. In particular, it is possible that the rate of the oscillations may change considerably around the center of the domain wall, i.e., in the region $|x|\lesssim w$. More remarkably, oscillations may appear only within an interval $|x|\lesssim x_{\text{osc}}$ if $v(x)^{2}+m(x)<0$ for $|x|<x_{\text{osc}}$ and $v(x)^{2}+m(x)=v_{L,R}^{2}+m_{L,R}\geq 0$ for $|x|>x_{\text{osc}}$ or conversely, only for $|x|>x_{\text{osc}}$ if $v(x)^{2}+m(x)\geq 0$ for $|x|<x_{\text{osc}}$ and $v(x)^{2}+m(x)=v_{L,R}^{2}+m_{L,R}<0$ for $|x|>x_{\text{osc}}$.
@@
Figure 2 show zero modes on the interval $-\infty<x<\infty$ in the case where the width of the smooth domain wall $w$ is smaller than or comparable with the characteristic decay lengths $\xi$, for different choices of the parameters $m_{0,1,2}$ and $v_{1,2}$, plotted using Eqs. 5 and 7. These figures show examples of modes with exponential decay or exponentially damped oscillations. The exponential decay of the modes is clearly visible, with the exception of a small region for $|x|<w<\xi$: the modes have "short-hair", i.e., they are nonfeatureless for $|x|\lesssim w$ but can be regarded as featureless for distances $|x|\gtrsim w$. Figure 2 show cases analogous to Fig. 2, but where the width of the smooth domain wall $w$ is larger than the characteristic decay lengths $\xi$. The most noticeable difference with the previous case is that the exponential decay of the modes is now visible only at large distances $|x|\gtrsim w>\xi$: the modes have "long-hair", i.e., they are nonfeatureless at finite length scales.
The local features of the zero modes become prominent when $w\gg\xi$ (long-hair), especially when the field $m(x)$ shows large variations around the domain wall, i.e., when the component $m_{2}$ of the field increases. This is particularly visible in Fig. 2(d) where $w>\xi$, with oscillations visible in a wide region close to the center where $v(x)^{2}+m(x)<0$ but disappear at larger distances where $v(x)^{2}+m(x)>0$, and to some extent in Fig. 2(d) where $w<\xi$, with oscillations visible in a narrower region. Conversely, in Fig. 2(f), oscillations seem to occur only at large distances where $v(x)^{2}+m(x)<0$ but disappear at smaller distances where $v(x)^{2}+m(x)>0$. Even more remarkably, the zero mode in Fig. 2(d) is localized far away from the center of the domain wall in correspondence of a point $|x_{1}^{*}|>\xi$, while the two zero modes in Fig. 2(f) are localized around the two points $|x_{1,2}^{*}|>\xi$ where the local topological invariant $W(x)$ changes and the field $m(x)$ changes its sign.
@@
$\widetilde{\varphi}(x)=\left(2\cosh\left(\frac{x}{2w}\right)\right)^{swv_{1}}\mathrm{e}^{s(2v_{0}+v_{1})x/2}\varphi(x),$ (21)
@@
$\widetilde{\varphi}_{1,2}(x)=$ $\frac{\mathrm{e}^{\pm({q_{L}}+{q_{R}})x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{{\pm w(q_{L}}-{q_{R}})}}F_{1,2}(x),$ (23)
@@
1) general case $\widetilde{\varphi}_{1,2}(x)=\dfrac{\mathrm{e}^{\pm\left(q_{L}+q_{R}\right)x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{\pm w\left(q_{L}-q_{R}\right)}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
Notice that we required different kinds of boundary conditions for the zero-energy solutions of the generalized Jackiw-Rebbi equation and for the associated Schrödinger equation. In the case of the generalized Jackiw-Rebbi equation, we are concerned about normalizable modes that satisfy the Dirichlet boundary conditions $\varphi(\pm\infty)=0$. These normalizable modes do not exist, in general, for the associated Schrödinger equation. In this case, we only require that the modes are bounded $|\widetilde{\varphi}(x)|<\infty$ for $x\to\pm\infty$. At long distances, these modes behave either as plane waves or decay exponentially.
@@
The bounded modes $\widetilde{\varphi}_{1,2}(x)$ on the infinite interval $-\infty<x<\infty$ with $\widetilde{W}_{L}=\widetilde{W}_{R}=1$, i.e., $E\geq V_{L,R}$, describe the transmission and reflection of particles with energies greater than the height of the domain wall $E\geq\max(V_{L,R})$. These modes are wavelike modes with oscillatory having momenta $\kappa_{L,R}=\mathrm{Im}(q_{L,R})=\sqrt{E-V_{L,R}}$, respectively on each side of the domain wall. Hence, these modes are completely delocalized. When $w\to 0$ and $V(x)=V_{L}=V_{R}$, this simply reduces to the case of a free particle with momentum $\kappa=\sqrt{E-V}$.
@@
Moreover, the bounded mode on the semi-infinite interval $0\leq x<\infty$ with $\widetilde{W}_{R}=1$, i.e., $E\geq V_{R}$ describes the reflection of a particle with momentum $\kappa=\sqrt{E-V_{R}}$ off an infinite wall with smooth spatial variations of the potential near $x=0$, which for $w\to 0$ reduces to the well-known case describing the reflection of a free particle scattering off an infinite wall.
@@
The bounded mode $\widetilde{\varphi}_{1}(x)$ with $\widetilde{W}_{L}=0$ and $\widetilde{W}_{R}=1$ with $V_{L}>E\geq V_{R}$ correspond via the duality transformation to the zero mode $\varphi_{1}^{s}(x)$ with $W_{L}=0$ and $|W_{R}|=1$ if we assume that $|m_{L}|>v_{L}^{2}$ and $v_{R}\neq 0$. On the right side, the bounded mode is oscillating and the zero node is exponentially decaying with damped oscillations and decay rate $|v_{R}|$. The oscillations have same momentum $\kappa_{R}=\mathrm{Im}(\sqrt{v_{R}^{2}+m_{R}})=\mathrm{Im}(\sqrt{V_{R}-E})$. Again, in the limiting cases where this momentum vanishes (i.e., if $E=V_{R}$) the bounded mode converges to a constant and the zero mode decays exponentially without oscillations. On the left side, the bounded mode and zero mode decay exponentially with decay rates $\mu_{L}=q_{L}$ and $\mu_{L}^{+}=-sv_{L}\pm q_{L}$, respectively. Same considerations can be made for the bounded mode $\widetilde{\varphi}_{2}(x)$ and zero mode $\varphi_{2}^{s}(x)$.
Finally, the bounded mode on the semi-infinite interval $0\leq x<\infty$ with $\widetilde{W}_{R}=1$, i.e., $E\geq V_{R}$, corresponds via the duality transformation to the zero mode if we assume $v_{R}\neq 0$. The bounded mode is oscillating and the zero node is exponentially decaying with damped oscillations, with same momentum $\kappa_{R}=\mathrm{Im}(\sqrt{v_{R}^{2}+m_{R}})=\mathrm{Im}(\sqrt{V_{R}-E})$. In the limiting case, $E=V_{R}$) the bounded mode converges to a constant and the zero node decays exponentially without oscillations. The duality between the bounded modes and the zero modes is summarized in Table 5.
@@
2 wavelike modes $\widetilde{\varphi}_{1,2}$ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ 2 zero modes $\varphi_{1,2}^{s}$
$V_{L}=E\Rightarrow$ const. at $\infty$- ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{L}^{2}+m_{L}=0\Rightarrow$ exp. decay at $\infty$- ∞
$V_{L}<E\Rightarrow$ oscill. at $\infty$- ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{L}^{2}+m_{L}<0\Rightarrow$ exp. damped oscill. at $\infty$- ∞
$V_{R}=E\Rightarrow$ const. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}=0\Rightarrow$ exp. decay at +$\infty$+ ∞
$V_{R}<E\Rightarrow$ oscill. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}<0\Rightarrow$ exp. damped oscill. at +$\infty$+ ∞
@@
1 Schokley-like mode $\widetilde{\varphi}_{1}$ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ 1 zero mode $\varphi_{1}^{s}$
exp. decay at $\infty$- ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ exp. decay at $\infty$- ∞
$V_{R}=E\Rightarrow$ const. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}=0\Rightarrow$ exp. decay at +$\infty$+ ∞
$V_{R}<E\Rightarrow$ oscill. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}<0\Rightarrow$ exp. damped oscill. at +$\infty$+ ∞